OpenStudy (anonymous):

Which set of three numbers could represent the lengths of the sides of a right triangle? 7, 9, 11 8, 15, 17 10, 15, 20 12, 18, 22

jhonyy9 (jhonyy9):

just you need try calculing 7^2 +9^2 = 11^2 or 8^2 +15^2 = 17^2 or 10^2 +15^2 = 20^2 or 12^2 +18^2 = 22^2 and will see in what case you get that the left side is equal right side hope so much that this will help you good luck bye